1.     Is the function 𝑓: ℤ → ℤ 𝑓(𝑥) = 𝑥 2 + 3 injective, surjective or bijective? Prove your assertions 


1
Expert's answer
2021-08-31T17:15:17-0400

Consider the function "f:\Z\to\Z,\ f(x)=x^2+3." Since for "x_1=-1" and "x_2=1\ne x_1" we have that "f(x_1)=(-1)^2+3=4=1^2+3=f(x_2)," we conclude that the function "f" is not injective. Taking into account that for "y=0\in\Z" the equation "x^2+3=0" has no integer roots, we conclude that "f^{-1}(0)=\emptyset," and hence the function "f" is not surjective. Therefore, the function "f" is not bijective.